### v1

@@ 1 Adversarial Attacks: Categories and Aims
Projected Gradient Descent (PGD) IFGSM attack represents one of the most established and efficient methods.It is a variant of Iterative FGSM, differing in their initialization strategy. More specifically, PGD initializes the example randomly within the sphere of interest determined by the $L_{\infty}$ norm and does random restarts, whereas IFGSM starts from the original point.
@@
Jacobian-based Saliency Maps JSM explore the forward derivatives to calculate the model gradients and discover which input regions/pixels that contribute the most to the predicted output. In fact, it uses saliency maps that highlight the most relevant or influential features within the input data concerning the model’s decision-making process. Hence, the adversary consists of targeting the perturbations of these critical input regions to create effective adversarial examples.
DeepFool deepfool is an iterative attack that aims to find the minimal perturbation required to misclassify an input by exploiting the model’s decision boundaries. It works by iteratively adjusting the input in the direction orthogonal to the decision boundary hyperplane, and it stops when it reaches an adversarial perturbation that causes a change in the model’s output classification. This approach often results in the smallest perturbations that are imperceptible but effective in causing misclassification.
AdvGAN AdvGAN uses Generative Adversarial Networks (GAN) GAN to create adversarial examples that not only have high attack success rate, but they are statistically distinguishable from the original ones. Indeed, AdvGAN operates iteratively, with a generator and a discriminator, engaged in a GAN-style training process. The generator aims to create perturbations that deceive the target model, while the discriminator learns to differentiate between original and perturbed instances. Through this adversarial training, AdvGAN seeks to produce adversarial examples that are both effective in fooling the target model and visually realistic, making them challenging for the target model to detect while maintaining a high attack success rate.
@@
Black-box adversarial attacks are less effective than white-box approaches because they often require a higher number of queries, but they can expose different adversarial examples and are better representative of external attack system simulations. Narodytska et al. narodytska2017simple performed a local-search-based attack. Chen et al. chen2017zoo and Bhagoji et al. bhagoji2018practical proposed black-box attacks based on gradient estimation spall2005introduction . Moon et al. moon2019parsimonious leveraged algorithms in combinatorial optimization. As opposed to white box gradient-based attacks, certain black-box adversarial attacks qiu2021black leverage Evolutionary Strategies to estimate the gradients used to perform a PGD-like attack. Alzantot et al. alzantot2019genattack recently reported about GenAttack, a gradient-free optimizer that uses Genetic Algorithms (GA) to produce subtle perturbations, which successfully fool state-of-the-art image recognition models with significantly fewer queries. Feature-Guided Black-Box (FGBB)s qiu2021black uses Scale Invariant Feature Transform to extract image features, emphasizing pixels that impact human visual perception. Then, crafting adversarial examples is framed as a strategic two-player game: one player manipulates features to minimize the distance to an adversarial example, while the other player adopts diverse strategies, i.e., cooperative, adversarial, or random. Using Monte Carlo tree search, FGBB systematically explores game states to uncover adversarial examples.
@@
In a broader sense, Meta Gradient Adversarial Attack MGAA is an innovative architecture designed to enhance cross-model transferability within gradient-based attack methods. This method operates through multiple iterations, where models are sampled from a model zoo in each iteration to create adversarial perturbations using the chosen model. These perturbations are then incorporated with previously generated ones. By leveraging multiple models, this approach simulates both white- and black-box settings, enhancing the effectiveness of the attacks.
The revelation of numerous inputs causing incorrect predictions by ML models, despite being anticipated as accurate, has garnered significant attention ren2020adversarial . This discovery sparked a race to create finely-tuned adversarial attacks, aiming to generate imperceptible noise capable of manipulating deep neural networks (DNNs) to produce erroneous outputs. However, criticism has surfaced regarding the practicality of these digital adversarial inputs in real-world scenarios. For example, even if an attacker possesses access to the autonomous driving car’s model, applying optimized imperceptible noise to a traffic sign on the road for misleading passing cars remains impractical. Although digital adversarial attacks may be successful in lab experiments or with models accessed through APIs, they are limited in the physical world. Consequently, physical adversarial attacks have emerged to execute unrestricted input alterations, which assemble all the synthetically-generated inputs without any $l_{p}$ norm bounding, but preserve the semantic identity of the source input. As a result, physical AXs are effective in complex real-world scenarios.
@@
Research into the causes of brittleness against adversarial attacks leads to the association-based statistical learning nature of supervised ML buckner2020understanding ; ilyas2019adversarial (Buckner, 2020; Ilyas et al., 2019). Nevertheless, their prevalence in modern ML models like deep neural networks is due to overparameterization that make them tend to exploit all patterns inherent in the data that contain predictive information, including such patterns that are inscrutable to human cognition or only associative but not causal for the prediction target freiesleben2022intriguing . This inherent tendency might clarify why various models trained on the same dataset can be deceived by identical adversarial examples.
@@ 2 Model-Centric Optimization Methods
Ensemble Adversarial Training tramer2017ensemble is proposed to mitigate the overfitting problem observed with fast single-step adversarial methods on large datasets. Indeed, this overfitting leads to a poor convergence of the global minimum, i.e., the obtained models often remain vulnerable to black-box attacks. Therefore, Ensemble Adversarial Training leverages existing static pre-trained models to craft adversarial examples required by AT, preventing the adversarially-trained model from weakening the credibility of adversarial examples.
Shared Adversarial Training mummadi2019defending is an extended version of AT that aims to model robustness against universal perturbations. It computes the loss gradients w.r.t a mini-batch of training inputs as a stack, then, a shared perturbation that is applied to the entire stack is derived. Afterward, these perturbations are iteratively added and clipped to meet predefined magnitude constraints that are required for adversarial training.
Due to its reliance on gradients, AT can be costly when creating adversarial examples, particularly for large datasets such as ImageNet. For this reason, Free Adversarial Training shafahi2019adversarial exploits the gradient information used in updating model parameters to craft the adversarial examples, thereby reducing computational costs.
In most cases, AT often leads to unfounded increases in the margin along decision boundaries, which negatively affects the original accuracy. The following strategies propose mitigations for better trade-offs between original and robust accuracy. Misclassification Aware Adversarial Training wang2019improving suggests the differentiation between the misclassified and correctly classified original inputs during adversarial training because the minimization step is more important for misclassified examples than the maximization one that is negligible. Friendly Adversarial Training zhang2020attacks targets the least adversarial examples (referred to as “friendly”) among the adversarial data that is confidently misclassified. When executing the adversarial training, the friendly adversarial examples can be identified by PGD attacks with an early stopping condition. Geometry-aware Instance-reweighted Adversarial Training zhang2020geometry proposes to leverage standard adversarial training with instance-level weights based on how difficult it is to attack an original data point. Indeed, the original data points that are close to the class boundary are less robust, which is why larger weights are assigned to them. Helper-based Adversarial Training rade2021helper introduces an additional set of wrongly-labeled examples during adversarial training, achieving a balanced trade-off between accuracy and robustness.
@@
The advantages of randomized smoothing cohen2019certified are: (i) It provides theoretical guarantees for robustness of models within certain probabilistic bounds, which is defined as certified robustness that guarantees a stable prediction for any input within a certain range. Thus, it offers a more principled approach to ensuring model robustness compared to empirical adversarial training strategies. (ii) It allows leveraging sophisticated pre-trained models as foundation to build robust ones through smoothing their decision boundaries using relevant datasets.
Fast Adversarial Training (FAT) chen2022efficient combines the strengths of randomized smoothing and single-step robust training methods, extracting the beneficial aspects from both approaches. On one hand, FAT performs randomized smoothing to optimize the inner maximization problem efficiently. On the other hand, it proposes a new initialization strategy, named backward smoothing, to improve the stability and robustness of a model using single-step robust training methods. This combination shows its effectiveness in improving model resilience against adversarial attacks while minimizing the risk of overfitting on adversary noises.
@@
Concerning robustness, a variety of loss functions have been used to incorporate specific objectives: triplet loss mao2019metric , minimising distance between true and false classes because adversarial attacks shift the internal representation towards the “false” class, consistency across data augmentation strategies tack2022consistency , and adding maximal class separation constraints mustafa2020deeply . In a similar manner, adapted regularisation can be designed to make models more robust. Li and Zhang li2021improved propose a PAC-Bayesian approach to tackle the memorization of training labels in fine-tuning. Chan et al. chan2019jacobian develop an approach to optimize the saliency of classifiers’ Jacobian by adversarially regularizing the model’s Jacobian in line with natural training inputs. Curvature Regularization moosavi2019robustness minimizes the curvature of the loss surface, which induces a more ”natural” behavior of the network because robustness requires quasi-linear behavior in the proximity of data points. Adversarial Weight Perturbation wu2020adversarial performs a double-perturbation mechanism that disturbs both inputs and weights to regularize the flatness of weight loss landscape and robustness gap.
@@
Another defense class of techniques against adversarial examples is Defensive Distillation papernot2016distillation , which trains a Distilled Network on the predictions from a previously trained neural network (the Initial Network). By replacing dataset labels with continuous prediction values, the Distilled Network is likely to be more resilient to adversarial attacks. In light of the promising results achieved by defensive distillation, Papernot and McDaniel papernot2017extending suggest combining the original label with the Initial model uncertainty when training the distilled network. Furthermore, Robust Soft Label Adversarial Distillation method is proposed to train robust small student DNNs using robust soft labels produced by an adversarially-trained larger teacher DNN to supervise the student training on natural and adversarial examples. In this method, the student DNN does not access the original complex labels through the training process. Low Temperature Distillation (LTD) chen2021ltd uses a relatively low temperature in the teacher model and different, but fixed, temperatures for the teacher and student models. Indeed, the temperature is a scaling parameter used to adjust the sharpness of the probability distribution over classes. Lower temperatures make the distribution sharper, emphasizing the most probable class, while higher temperatures result in a softer, more uniform distribution over classes. LTD lowers the temperature of the teacher model used for crafting soft labels to boost the student model’s robustness by only lowering the temperature of the teacher model for crafting soft labels without encountering the gradient masking problem (occurs when the loss gradients w.r.t inputs significantly diminish during the training of the distilled model).
@@ 3 Post-training Model Enhancements
The strategies, which involve changing training data or updating the model, require retraining or fine-tuning, which is costly since modern neural networks are complex. Moreover, pretrained models can sometimes be obtained from a third party or training data may be confidential, making access to the whole in-data distribution challenging. Importantly, there is a disconnection between the assessment techniques and the improvement processes for model robustness. There is no guarantee that the new optimized or upgraded model will respond correctly to the counterexamples discovered during the assessment using most amelioration techniques. Robust models are often evaluated on adversarial examples or shifted inputs generated by assessment methods. It is important to show that the amelioration strategy does not overfit on counterexamples, but it is often overlooked how successful it is at fixing the revealed erroneous behaviors. Therefore, researchers have proposed post-training, model-level repair of DNNs, i.e., repair through the modification of the weights of an already trained model. In conventional software systems, these narrowly-defined modifications are called patches. When it comes to deep learning, the patch aims to fix the model’s erroneous behaviors against a specific subset of counterexamples. Although the high complexity of modern ML models makes it difficult to perform provable model repairs (i.e., the counterexample provided will always be corrected), heuristic repair strategies that do not guarantee to repair all observed counterexamples still have benefits when compared to retraining or fine-tuning the model. Indeed, the original neural network likely went through training and functions properly on most inputs, so applying small, less impactful but precise changes to fix brittleness makes sense. CAusality-based REpair (CARE) sun2022causality identifies the “‘guilty” neurons (i.e., the ones that caused the false prediction) using causality-based fault localization, then, modifies the weights of these identified neurons to reduce the misbehaviors. Arachne sohn2023arachne is similar to CARE, but ensures no disturbance of the model’s correct behaviors. Indeed, it uses differential evolution to generate effective patches for the localised weights that can fix specific mispredictions of a DNN without drastically reducing its clean accuracy. GenMuNN wu2022genmunn ranks the weights based on their influences on the model’s predictions. It then generates mutants using the computed ranks and evolves them using genetic algorithms to increase the chances of finding mutants that satisfy the stopping criteria. By tracking the training history, NeuRecover tokui2022neurecover finds weights that have changed significantly over time. A weight becomes a subject for repair when it is no longer contributing to correct predictions in the earlier stage of training but is leading to incorrectly-predicted inputs. Similarly to NeuRecover, I-Repair henriksen2022repairing modifies localised weights to influence predictions for a specific set of fault-revealing inputs while minimising the impact on correct predictions. NNrepair [68] uses fault localisation to identify suspicious weights, and then uses constraint solving to modify them marginally. There are provable model repairs like PRDNN sotoudeh2021provable , REASSURE fu2021sound , and Minimal Modifications of DNNs (MMDNN) goldberger2020minimal . However, these approaches are not scalable to large DNNs, only support ReLU activation, and often do not support polytope repair or multilayer repairs.

### v2

@@ 1 Adversarial Attacks: Categories and Aims
Projected Gradient Descent (PGD) PGD attack represents one of the most established and efficient methods.It is a variant of Iterative FGSM, differing in their initialization strategy. More specifically, PGD initializes the example randomly within the sphere of interest determined by the $L_{\infty}$ norm and does random restarts, whereas IFGSM starts from the original point.
@@
Papernot et al. JSM introduced a jacobian-based class of attacks to craft adversarial examples based on computing forward derivatives. Knowing the network architecture, an adversary can construct saliency maps that identify features of the input that most significantly impact output classification. Indeed, saliency maps simonyan2013deep were originally designed for visually explaining the input features the classifier used to generate its output decisions. The map indicates how much influence each input feature (e.g. pixel) has on predicting a specific class. By increasing a few high-saliency pixels, $x(i)$, an attacker can influence the model to predict the adversarial class, $t$, that does not match the true class label, $y$, of a given input $x$, and this may result in a misclassification of the input.
DeepFool deepfool algorithm iterates to find the smallest perturbation that could misclassify an input. It represents the classifier’s decision space with linear hyperplane boundaries and iteratively shifts the image towards the closest boundary until misclassification. The algorithm takes an input $x$ and a classifier $f$, outputs the minimal perturbation, initializes the adversarial input with the original, and loops while the true label matches the adversarial label. It calculates the minimal perturbation by projecting the input onto the closest hyperplane, adds it to the input, and tests for misclassification. Finally, it returns the minimal perturbation.
Xiao et al. AdvGAN propose to generate adversarial examples with generative adversarial networks (GANs), which can learn and approximate the distribution of original instances. Using GAN-style learning, they train the generator to create perturbations that deceive the target model, while the discriminator is optimized to differentiate between original and perturbed instances. Once the generator is trained, it can generate adversarial perturbations efficiently for any instance to derive a visually-realistic crafted version that will likely fool the target model.
@@
Black-box adversarial attacks are less effective than white-box approaches because they often require a higher number of queries, but they can expose different adversarial examples and are better representative of external attack system simulations. Narodytska et al. narodytska2017simple performed a local-search-based attack. Chen et al. chen2017zoo and Bhagoji et al. bhagoji2018practical proposed black-box attacks based on gradient estimation spall2005introduction . Moon et al. moon2019parsimonious leveraged algorithms in combinatorial optimization. As opposed to white box gradient-based attacks, certain black-box adversarial attacks qiu2021black leverage Evolutionary Strategies to estimate the gradients used to perform a PGD-like attack. Alzantot et al. alzantot2019genattack recently reported about GenAttack, a gradient-free optimizer that uses Genetic Algorithms (GA) to produce subtle perturbations, which successfully fool state-of-the-art image recognition models with significantly fewer queries. Lu et al. FGBB propose to employ object detection techniques such as SIFT (Scale Invariant Feature Transform) to extract features from an image. The features are converted into a saliency distribution, where pixels that affect the composition of the image are given high probabilities. Then, crafting adversarial examples is framed as a strategic two-player game: one player manipulates features to minimize the distance to an adversarial example, while the other player adopts diverse strategies, i.e., cooperative, adversarial, or random. The resulting game states can be explored efficiently by Monte Carlo tree search to produce adversarial examples.
@@
Inspired by meta-learning, Yuan et al. MGAA introduced a versatile method designed to enhance cross-model transferability within gradient-based attack frameworks. Their approach seamlessly integrates with any existing attack method and focuses on narrowing the disparity between white-box and black-box attack gradients. By randomly selecting models from a model zoo and iteratively conducting both white-box and black-box attacks on diverse tasks, they effectively enhance the transferability of adversarial examples across models in black-box scenarios.
The revelation of numerous inputs causing incorrect predictions by ML models, despite being anticipated as accurate, has garnered significant attention ren2020adversarial . This discovery sparked a race to create finely-tuned adversarial attacks, aiming to generate imperceptible noise capable of manipulating deep neural networks (DNNs) to produce erroneous outputs. For more comprehensive details, we refer to the survey by Costa et al. costa2023deep , which delves into adversarial attacks, including methodologies and recent advancements. As the adversaries grow in popularity, criticism has surfaced regarding the practicality of these digital adversarial inputs in real-world scenarios. For example, even if an attacker possesses access to the autonomous driving car’s model, applying optimized imperceptible noise to a traffic sign on the road for misleading passing cars remains impractical. Although digital adversarial attacks may be successful in lab experiments or with models accessed through APIs, they are limited in the physical world. Consequently, physical adversarial attacks have emerged to execute unrestricted input alterations, which assemble all the synthetically-generated inputs without any $l_{p}$ norm bounding, but preserve the semantic identity of the source input. As a result, physical AXs are effective in complex real-world scenarios.
@@
Research into the causes of brittleness against adversarial attacks leads to the association-based statistical learning nature of supervised ML buckner2020understanding ; ilyas2019adversarial . Nevertheless, their prevalence in modern ML models like deep neural networks is due to overparameterization that make them tend to exploit all patterns inherent in the data that contain predictive information, including such patterns that are inscrutable to human cognition or only associative but not causal for the prediction target freiesleben2022intriguing . This inherent tendency might clarify why various models trained on the same dataset can be deceived by identical adversarial examples.
@@ 2 Model-Centric Optimization Methods
Adversarial training with one-step attacks is prone to overfitting, especially with large datasets. The overfitting problem hinders the convergence to a global minimum, which makes the adversarially-trained models vulnerable to black-box attacks. Tramer et al. tramer2017ensemble leverage multiple pre-trained models to produce adversarial examples for AT. They aim to remove the tight coupling between the on-training model and the adversarial attack, which could discredit the adversarial examples in the long run. Mummadi et al. mummadi2019defending propose to compute the gradients w.r.t a batch of training inputs rather than one single input, then, the gradients are merged to create a shared perturbation for the entire batch of inputs. Once these shared perturbations have been iteratively refined and clipped, they are used to augment training data during AT.
The high cost of generating strong adversarial examples via gradients makes adversarial training impractical on large-scale problems like ImageNet. Shafahi et al. shafahi2019adversarial develop a cost-effective AT version that eliminates the overhead cost of generating adversarial examples by recycling the gradient information computed when updating model parameters. Indeed, both model parameters and image perturbations are updated simultaneously during the backward pass, rather than using separate gradient computations. This intuitive sharing of gradients has the same computational cost as non-adversarial training and can be 3-30 times faster than conventional adversarial training.
In most cases, AT often leads to unfounded increases in the margin along decision boundaries, which negatively affects the original accuracy. To achieve a better trade-off between original and robust accuracy, several mitigation strategies have been proposed. Wang et al. wang2019improving suggested the differentiation between the misclassified and correctly classified original inputs during adversarial training because the minimization step is more important for misclassified examples than the maximization one that is negligible. Zhang et al. zhang2020attacks proposed to selectively consider the least adversarial examples (referred to as “friendly”) among the adversarial data that is confidently misclassified. These “friendly” adversarial examples can be systematically identified by adding early-stopping criteria to iterative attack methods. Zhang et al. zhang2020geometry added weights for data points used in AT based on their potential to derive adversarial examples. Close to the class boundary, the original data points are more susceptible to being turned into adversarial examples, which is why larger weights are assigned to them. Rade et al. rade2021helper found that the AT induces an unexpected increase in the margin along certain adversarial directions, thereby degrading the original accuracy. Hence, they proposed to reduce this effect by incorporating additional incorrectly-labelled data during training, which indeed improves accuracy without compromising robustness. To find out more details, we recommend Costa et al.’s survey costa2023deep , which elaborates on the methodologies and advances of adversarial training.
@@
The advantages of randomized smoothing cohen2019certified are: (i) It provides theoretical guarantees for robustness of models within certain probabilistic bounds, which is defined as certified robustness that guarantees a stable prediction for any input within a certain range. Thus, it offers a more principled approach to ensuring model robustness compared to empirical adversarial training strategies. (ii) It allows leveraging sophisticated pre-trained models as foundation to build robust ones through smoothing their decision boundaries using relevant datasets. Fast Adversarial Training chen2022efficient combines the strengths of randomized smoothing and single-step adversarial training to make the most of both approaches. On one hand, Fast AT applies randomized smoothing for better optimization of the inner maximization problem. On the other hand, Fast AT proposes backward smoothing initialization over single-step random initialization, which enhances smoothing in the -perturbation ball.
@@
Concerning robustness, a variety of loss functions have been used to incorporate specific objectives: triplet loss mao2019metric , minimising distance between true and false classes because adversarial attacks shift the internal representation towards the “false” class, consistency across data augmentation strategies tack2022consistency , and adding maximal class separation constraints mustafa2020deeply . In a similar manner, adapted regularisation can be designed to make models more robust. Li and Zhang li2021improved propose a PAC-Bayesian approach to tackle the memorization of training labels in fine-tuning. Chan et al. chan2019jacobian develop an approach to optimize the saliency of classifiers’ Jacobian by adversarially regularizing the model’s Jacobian in line with natural training inputs. Moosavi et al. moosavi2019robustness studied the impact of AT on the curvature of the loss surface with respect to inputs. They found that AT reduced the loss curvature of the network, resulting in a more linear behavior. Hence, they proposed curvature regularization (CURE), which is a new regularizer that directly minimizes curvature of the loss surface to promote quasi-linear behavior in the vicinity of data points. CURE leads to adversarial robustness that is comparable to that of adversarial training. Furthermore, Wu et al. wu2020adversarial found a correlation between robust generalization and the flatness of weight loss landscapes resulting from the AT. Thus, they proposed Adversarial Weight Perturbation (AWP), which regularizes weight loss landscapes by adversarially perturbing both inputs and weights. By injecting systematically worst-case weight perturbations, AWP can be easily incorporated into any existing adversarial training approaches with little overhead.
@@
Another defense class of techniques against adversarial examples is Defensive Distillation papernot2016distillation , which trains a Distilled Network on the predictions from a previously trained neural network (the Initial Network). By replacing dataset labels with continuous prediction values, the Distilled Network is likely to be more resilient to adversarial attacks. In light of the promising results achieved by defensive distillation, Papernot and McDaniel papernot2017extending suggest combining the original label with the Initial model uncertainty when training the distilled network. Zi et al. zi2021revisiting proposed to train robust small student DNNs using the predictions produced by an adversarially-trained larger teacher DNN as robust soft labels. As a result, the student DNN is guided to learn both natural and adversarial examples in all loss terms. Low Temperature Distillation (LTD) chen2021ltd uses a relatively low temperature in the teacher model and different, but fixed, temperatures for the teacher and student models. Indeed, the temperature is a scaling parameter used to adjust the sharpness of the probability distribution over classes. Lower temperatures make the distribution sharper, emphasizing the most probable class, while higher temperatures result in a softer, more uniform distribution over classes. LTD lowers the temperature of the teacher model used for crafting soft labels to boost the student model’s robustness by only lowering the temperature of the teacher model for crafting soft labels without encountering the gradient masking problem (occurs when the loss gradients w.r.t inputs significantly diminish during the training of the distilled model).
@@ 3 Post-training Model Enhancements
The strategies, which involve changing training data or updating the model, require retraining or fine-tuning, which is costly since modern neural networks are complex. Moreover, pretrained models can sometimes be obtained from a third party or training data may be confidential, making access to the whole in-data distribution challenging. Importantly, there is a disconnection between the assessment techniques and the improvement processes for model robustness. There is no guarantee that the new optimized or upgraded model will respond correctly to the counterexamples discovered during the assessment using most amelioration techniques. Robust models are often evaluated on adversarial examples or shifted inputs generated by assessment methods. It is important to show that the amelioration strategy does not overfit on counterexamples, but it is often overlooked how successful it is at fixing the revealed erroneous behaviors. Therefore, researchers have proposed post-training, model-level repair of DNNs, i.e., repair through the modification of the weights of an already trained model. In conventional software systems, these narrowly-defined modifications are called patches. When it comes to deep learning, the patch aims to fix the model’s erroneous behaviors against a specific subset of counterexamples. Although the high complexity of modern ML models makes it difficult to perform provable model repairs (i.e., the counterexample provided will always be corrected), heuristic repair strategies that do not guarantee to repair all observed counterexamples still have benefits when compared to retraining or fine-tuning the model. Indeed, the original neural network likely went through training and functions properly on most inputs, so applying small, less impactful but precise changes to fix brittleness makes sense. CAusality-based REpair (CARE) sun2022causality identifies the “‘guilty” neurons (i.e., the ones that caused the false prediction) using causality-based fault localization, then, modifies the weights of these identified neurons to reduce the misbehaviors. Arachne sohn2023arachne is similar to CARE, but ensures no disturbance of the model’s correct behaviors. Indeed, it uses differential evolution to generate effective patches for the localised weights that can fix specific mispredictions of a DNN without drastically reducing its clean accuracy. GenMuNN wu2022genmunn ranks the weights based on their influences on the model’s predictions. It then generates mutants using the computed ranks and evolves them using genetic algorithms to increase the chances of finding mutants that satisfy the stopping criteria. By tracking the training history, NeuRecover tokui2022neurecover finds weights that have changed significantly over time. A weight becomes a subject for repair when it is no longer contributing to correct predictions in the earlier stage of training but is leading to incorrectly-predicted inputs. Similarly to NeuRecover, I-Repair henriksen2022repairing modifies localised weights to influence predictions for a specific set of fault-revealing inputs while minimising the impact on correct predictions. NNrepair NNRepair uses fault localisation to identify suspicious weights, and then uses constraint solving to modify them marginally. There are provable model repairs like PRDNN sotoudeh2021provable , REASSURE fu2021sound , and Minimal Modifications of DNNs (MMDNN) goldberger2020minimal . However, these approaches are not scalable to large DNNs, only support ReLU activation, and often do not support polytope repair or multilayer repairs.